OpenStudy (anonymous):

Which is the cost per square inch of a pizza with diameter 20 inches if the cost of the pizza is $16? Be sure to round your answer to the nearest cent.

OpenStudy (anonymous):

a. about 1 cent b. about 5 cents c. about twenty-five cents d. about eighty cents

OpenStudy (anonymous):

314.16 would be the area of pizza. So 1600/314.16 = 5.0929456269... so about 5 cents (B)

OpenStudy (anonymous):

A=πr^2 would be your formula for area of a circle.

OpenStudy (anonymous):

thank you!

OpenStudy (akashdeepdeb):

Cost of the pizza = $16 Now one thing we should know here is that, a pizza is priced on the basis of its size. So a small pizza may have a lesser price. Now, the diameter is: 20 inches. So basically the radius [which is half of a diameter is]: 10 inches. The area of a pizza [which is circular] would be given by the equation: \(\pi * r^2\) So if the \(r\) = 10 inches and \(\pi\) = 3.14 Then the area is 314 inches\(^2\) Now this is the area of the whole pizza. Price of the whole pizza = $16 Price per unit area = \(\frac{16}{314}$/inches^2\) Getting this? :)

OpenStudy (anonymous):

so b?

OpenStudy (akashdeepdeb):

Yes, in cents it should approximately be 'B'. Getting this concept behind it?